OpenStudy (anonymous):

A cone has a diameter of 12 in. and a height of 16 in. What is the volume of the cone? Use 3.14 as an approximation for straight pi and express your final answer in hundredths.

563blackghost (563blackghost):

Formula for cone... \[V = \frac{ 1 }{ 3 } \pi r^2 h\]

563blackghost (563blackghost):

how do we get radius?

OpenStudy (anonymous):

by dividing 12 by 2

563blackghost (563blackghost):

Correct! so it will be 6.... now we know what the height and the radius so we input it into the formula....\[V = \frac{ 1 }{ 3 } (3.14) 6^2 (16)\]

563blackghost (563blackghost):

@iamme219

OpenStudy (anonymous):

ok let me solve it

OpenStudy (anonymous):

i got 602.88 @563blackghost

563blackghost (563blackghost):

your answer is correct...